# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ IV.2.3 Z boson triphoton decays
As seen in Table 10, the $\rm Z\to 3\gamma$ decay is extremely suppressed in the SM ($\mathcal{B}=6.4\cdot 10^{-10}$, as computed here) making of such a final state a particularly clean means to search for exotic Z decays into BSM particles with EW couplings. Many new physics scenarios contain new light (pseudo)scalar or tensor particles that couple to the electroweak (or Higgs) bosons and decay primarily to photons. Searches for $\mathrm{Z}\to\gamma\,a(\gamma\gamma)$, with $a$ being an ALP [156, 20] or a massive graviton [21] decaying into photons, have attracted an increasing interest in the last years [19]. Figure 18 shows possible diagrams leading to the triphoton decay of the Z boson. The direct decay (left diagram) and the exclusive radiative decays into C-even mesons followed by their diphoton decays (center) can mimic the BSM processes signals (right), and constitute backgrounds to the latter. The direct and mesonic backgrounds have been neglected so far in all ALP and massive-graviton upper limits set in studies based on current and future triphoton Z decay data. Such an assumption is justified so far given the null sensitivity to such rare SM processes, but it will not be the case at the FCC-ee facility where, as we see below, a few thousands such events are expected during the whole Z-pole run. Here, we quantify the triphoton branching fractions from meson decays by combining the exclusive $\rm Z\to\gamma+M$ results for spin-0,2 mesons of Tables 1113 with their corresponding two-photon branching fractions [58]. The results are listed in Table 15. The sum of all exclusive mesonic decay channels amounts to $\mathcal{B}=1.8\cdot 10^{-10}$, representing an increase of about 30% from the direct $3\gamma$ decay. The yields of SM triphoton decays will therefore amount to about 1 000 events for the $N(\rm Z)=6\cdot 10^{12}$ bosons expected to be collected during the full Z-pole operation at the FCC-ee. Searches for ALPs and new spin-2 particles will have to be carried out on top of such a relatively large number of background events that have been neglected so far in the derivation of ALP limits at future $\mathrm{e}^{+}\mathrm{e}^{-}$ facilities [156].
@@
Table 15: Compilation of exclusive Z decays into triphoton final states. The predictions of relevant $\rm Z\to\gamma+M$ radiative decays listed in Tables 1113 are combined with the diphoton branching fractions of the produced mesons, $\rm M\to\gamma\gamma$ [58, 157].
@@
$Z$longrightarrow $\gamma\gamma\gamma$ $6.6$$\times~{}10^{-10}$
@@ IV.2.4 Z boson decays into a W boson plus a meson
The exclusive Z boson decays into $\rm W^{\mp}$ boson plus a charged meson $\rm M^{\pm}$ correspond to the diagrams of Fig. 15 with W bosons in the final state. Theoretical estimations for such decays have been given in Ref. [44] for light-quark and $\rm D^{\pm}$ mesons. We estimate here the rest of missing exclusive decays: $\rm Z\to W^{\mp}+\{D^{*\pm},D^{*\pm}_{s},B^{\pm},B^{*\pm},B^{*\pm}_{c},B^{\pm}_{c}\}$, using the expression
@@
$\rm Z\,\to\,W$ $+$ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-ee
$\rm Z\,\to\,W^{\mp}$ + $\rm\pi^{\pm}$  $(1.51\pm 0.01)$$\times~{}10^{-10}$ SCET+LCDA [44] $<$ $7.0$×$\times~{}10^{-5}$ [145]
$\rm\rho^{\pm}$  $(4.0\pm 0.1)$$\times~{}10^{-10}$ SCET+LCDA [44] $<$ $8.3$×$\times~{}10^{-5}$ [145]
$\rm K^{\pm}$  $(1.16\pm 0.01)$$\times~{}10^{-11}$ SCET+LCDA [44]
$\rm K^{*\pm}$  $(1.96\pm 0.12)$$\times~{}10^{-11}$ SCET+LCDA [44]
$\rm D^{\pm}$  $(1.99\pm 0.17)$$\times~{}10^{-11}$ SCET+LCDA [44]
$\rm D_{s}^{\pm}$  $(6.04\pm 0.30)$$\times~{}10^{-10}$ SCET+LCDA [44]
@@
Figure 19: Branching fractions (in negative log scale) of exclusive $\rm Z\to W^{\pm}+meson$ decays. The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-ee reach based only on the total number of Z bosons to be produced.
@@ IV.3 Radiative Z boson leptonium decays
Figure 15 shows the diagrams for the decay of the Z boson into a photon plus a leptonium state. Depending on the relative helicities of the outgoing leptons, their bound state can be in the spin singlet (paraleptonium, $(\ell\ell)_{0}$) or triplet (ortholeptonium, $(\ell\ell)_{1}$) states. No calculation for such processes exists to our knowledge, but the key ingredients can be obtained from the determination of the $\rm\mathrm{e}^{+}\mathrm{e}^{-}\to\gamma+(\tau^{+}\tau^{-})_{0}$ cross section at the Z pole provided in the ditauonium studies of Ref. [158]. The motivations for their study are the same as for the similar Higgs decays, namely, the search for LFV or LFUV phenomena.
@@
for the ortho-leptonium one, where $n$ is the principal quantum number of the leptonium resonance, and $s_{w}\equiv\sin\theta_{\mathrm{W}}$, and $c_{w}\equiv\cos\theta_{\mathrm{W}}$ are the sine and cosine of the Weinberg angle (Table 2). In both expressions above, $\alpha(0)$ and $\alpha(m_{\mathrm{Z}})$ are the electromagnetic coupling evaluated at zero [^10] The leptonium wavefunction is proportional to $\alpha(0)^{3}$, Eq. (24), which combined with the emission of a final-state onshell photon, gives the $\alpha(0)^{4}$ dependence shown in Eqs. (42) and (43). and at the Z pole mass, respectively, that, together with the rest of parameters, are listed in Table 2. The obtained branching fractions are tabulated in Table 17, and plotted in Fig. 21. The dependence of these Z-boson radiative rates on $\alpha^{6}\,m_{\ell\ell}^{2}$, leads to vanishingly small branching fractions, $\mathcal{O}(10^{-13}$$10^{-23})$, with ditauonium (positronium) featuring the largest (smallest) values. No experimental search has been conducted to date, although the relatively long lifetime of the leptonium objects (significantly boosted in the decay of the much heavier Z boson) would lead to a clean signature characterized by a displaced vertex from secondary decays of the $(\ell^{+}\ell^{-})$ into photons, $\rm e^{\pm}$, and/or $\mu^{\pm}$ [117], akin to many BSM long-lived particles (but with an invariant mass at the corresponding known leptonium mass). Given the negligible rates, only an FCC-ee run at the Z pole would be able to provide limits approaching the SM values for the $\rm Z\to\gamma+(\tau^{+}\tau^{-})_{1}$ case.
@@
$\rm Z\,\to\,\gamma$ $+$ $\rm(\ell\ell)$ Branching fraction Framework 2024 HL-LHC FCC-ee
$\rm Z\,\to\,\gamma$ + $\rm(ee)_{1}$  $7.3$$\times~{}10^{-21}$ (this work)
$\rm(\mu\mu)_{1}$  $3.1$$\times~{}10^{-16}$ (this work)
$\rm(\tau\tau)_{1}$  $8.9$$\times~{}10^{-14}$ (this work)
$\rm(ee)_{0}$  $4.7$$\times~{}10^{-23}$ (this work)
$\rm(\mu\mu)_{0}$  $2.0$$\times~{}10^{-18}$ (this work)
$\rm(\tau\tau)_{0}$  $5.7$$\times~{}10^{-16}$ (this work)
Figure 21: Branching ratios (in negative log scale) of exclusive $\rm Z\to\gamma+leptonium$ decays. Theoretical predictions computed here are shown as blue bars. The red vertical line indicates the expected FCC-ee reach based on the total number of Z bosons to be produced.
@@ IV.4 Exclusive Z boson decays into two mesons
Figure 22 displays the diagrams of exclusive Z boson decays into two mesons, with contributions from direct quark decays (left and center), as well as from indirect $\rm V^{*}\to M$ transitions (right). Due to the coincidence of diagrams with those of the Higgs boson (Fig. 11), and since the Z boson yields at colliders are $10^{3}$$10^{6}$ times larger (Table 1), the study of such processes provides valuable information for the corresponding Higgs rates. The two-body decays into the same pair of pseudoscalar mesons, $\rm Z\to M+M$, are quantum-mechanically forbidden due to the presence of two identical final-state particles and conservation of angular momentum (thus violating the spin-statistics theorem), but the mesons M can be in a vector state. The decays of Z bosons into double quarkonia, first studied in 1990 [159], have been investigated as a means to provide clean information on the quarkonium bound-state dynamics at large momentum transfers. As explained below, there is still a relatively large uncertainty in the theoretical predictions, but this is still the most promising place to search for exclusive double-charmonia/bottomonia decays.
@@
To our knowledge, there are no calculations of exclusive decays of the Z boson involving light vector mesons, but only into charmonium and/or bottomonium states. We have evaluated here for the first time the double light vector meson decay $\rm Z\to VM+VM$ via the two direct (left and center) diagrams of Fig. 22, using the EFT + NRQM formalism of [160] in which the width can be written as
@@
for $N_{\mathrm{c}}=3$ colours, and without higher-order QCD corrections. We can only estimate the rate for the $\rm Z\to\phi+\phi$ channel because it is the only pure $|\mathrm{q}\overline{\mathrm{q}}\rangle$ state among the light mesons. Using Eqs. (44)–(47) and the parameters of Table 2, we obtain a $\mathcal{B}=2.1\cdot 10^{-12}$ rate (Table 18). There is no $\rm Z\to\phi+\phi$ experimental search performed to date, and only a machine like FCC-ee can provide enough Z bosons to start producing the decay. The table also quotes the forbidden $\rm Z\to\pi^{0}+\pi^{0}$ channel, for which the CDF upper bound could be improved by a factor of 100 at the HL-LHC. To provide a reasonable evaluation for exclusive double decays into other light mesons, rotation from flavour eigenstates $|\mathrm{u}\overline{\mathrm{u}}\rangle,|\mathrm{d}\overline{\mathrm{d}}\rangle$ to physical eigenstates such as $\rho,\omega,\eta...$ would need to be performed. One can, however, anticipate branching fractions of the same order as the double-$\phi$ channel, $\mathcal{O}(10^{-12})$, since they only differ by a rotation and they are enhanced by the second and third diagram of Fig. 22 involving extra $\rm\gamma^{*},W^{*}$ contributions. The scenario where Z decays into light pseudoscalar plus vector mesons suffers from the same complications of mixed states, but their rates are expected to be of the same order-of-magnitude as the double vector meson case because they have the same enhancing effect from the photon propagator contribution (Fig. 22, center) [160].
@@
$Z\,\to$ M $+$ M Branching fraction Framework 2024 HL-LHC at FCC-ee
$\rm Z\,\to$ $\rm\pi^{0}$ + $\rm\pi^{0}$ (forbidden) $<$ $1.5$×$\times~{}10^{-5}$ [138] $\lesssim 2.1\times~{}10^{-7}$
$\rm\pi^{+}$ + $\rm\pi^{-}$  $1.7$$\times~{}10^{-12}$ SCET+LCDA [161]
$\rm K^{+}$ + $\rm K^{-}$  $8.3$$\times~{}10^{-13}$ SCET+LCDA [161]
@@
There is a recent calculation of exclusive Z decays into a pair of charged mesons, $\rm Z\to\pi^{+}\pi^{-},K^{+}+K^{-}$ [161]. Also, the branching fractions $\rm Z\to B_{c}^{(*)+}+B_{c}^{(*)-}$ can be estimated to be in the $\mathcal{O}(10^{-8})-\mathcal{O}(10^{-11})$ range [162, 163, 164], where the variation of results within a factor of 20 depends on the different potential models [163] or initial quark masses [164] used.
Tables 19 and 20 list the theoretical predictions and experimental limits for concrete two-meson decay modes with pairs of charmonium mesons and bottomonium plus other mesons, respectively. The same results are shown in graphical form in Figs. 23 and 24. The decay rates are very small, $\mathcal{B}(\rm Z\to M+M)\lesssim 10^{-10}$, and quite sensitive to the model details and ingredients. All heavy-quarkonium channels have rates obtained within the NRQCD + LDMEs formalism. Alternative LC calculations show a good agreement with the latter [165] but, unfortunately, they can only provide predictions for final states with different mesons or with mesons in different excited states. The NRQCD predictions from the work [165] show discrepancies with other later works, because they did not consider diagrams involving a photon propagator (Fig. 22, center) that can enhance the rates by more than two orders-of-magnitude. In contrast, this latter study provides results for some excited states that have not been calculated elsewhere (although their results for channels such as $\rm Z\to\chi_{c1}+\chi_{c1},\,\chi_{c2}+\chi_{c2},\,\chi_{c0}+\chi_{c2}$,…, which cannot be predicted using LC formalism, may need to be revisited for the same reasons stated above). In addition, recent calculations including higher-order QED [166] and QCD + QED [167] corrections yield significant enhancements in the branching ratios for double-quarkonia production. All that said, and within the relatively large theoretical uncertainties for those processes, the largest branching fractions expected are for double-charmonium with $\mathcal{O}(10^{-10})$ rates. Despite a few searches carried out at LEP (and one at the LHC), experimental observation will only be possible at FCC-ee for about half of them.
@@
$Z\,\to$ M $+$ M Branching fraction Framework 2024 HL-LHC at FCC-ee
$\rm Z\,\to$ $\rm\eta_{c}$ + $\rm J/\psi$  $(1.5\pm 0.4)$$\times~{}10^{-11}$ NRQCD/NRCSM [160]
@@
 $2.7$$\times~{}10^{-14}$ NRQCD+LDME [165]
$\rm\chi_{c0}$  $2.3$$\times~{}10^{-12}$ NRQCD+LDME [165]
 $(2.3\pm 1.0)$$\times~{}10^{-12}$ LC+LCDA [165]
$\rm\chi_{c1}$  $5.4$$\times~{}10^{-14}$ NRQCD+LDME [165]
$\rm h_{c}$  $2.1$$\times~{}10^{-13}$ NRQCD+LDME [165]
 $(1.0\pm 0.5)$$\times~{}10^{-12}$ LC+LCDA [165]
$\rm\chi_{c2}$  $9.7$$\times~{}10^{-13}$ NRQCD+LDME [165]
 $(4.6\pm 2.0)$$\times~{}10^{-12}$ LC+LCDA [165]
$\rm J/\psi$ + $\rm J/\psi$  $(1.1\pm 0.3)$$\times~{}10^{-10}$ NRQCD/NRCSM [160] $<$ $1.1$×$\times~{}10^{-6}$ [107] $\lesssim 1.7\times~{}10^{-7}$
 $\left(1.11^{+0.34}_{-0.24}\right)$$\times~{}10^{-10}$ NRQCD+LDME [167]

### v4

@@ IV.2.3 Z boson triphoton decays
As seen in Table 10, the $\rm Z\to 3\gamma$ decay is extremely suppressed in the SM ($\mathcal{B}=6.4\cdot 10^{-10}$, as computed here) making of such a final state a particularly clean means to search for exotic Z decays into BSM particles with EW couplings. Many new physics scenarios contain new light (pseudo)scalar or tensor particles that couple to the electroweak (or Higgs) bosons and decay primarily to photons. Searches for $\mathrm{Z}\to\gamma\,a(\gamma\gamma)$, with $a$ being an ALP [Bauer:2018uxu, dEnterria:2021ljz] or a massive graviton [dEnterria:2023npy] decaying into photons, have attracted an increasing interest in the last years [Agrawal:2021dbo]. Figure 18 shows possible diagrams leading to the triphoton decay of the Z boson. The direct decay (left diagram) and the exclusive radiative decays into C-even mesons followed by their diphoton decays (center) can mimic the BSM processes signals (right), and constitute backgrounds to the latter. The direct and mesonic backgrounds have been neglected so far in all ALP and massive-graviton upper limits set in studies based on current and future triphoton Z decay data. Such an assumption is justified so far given the null sensitivity to such rare SM processes, but it will not be the case at the FCC-ee facility where, as we see below, a few thousands such events are expected during the whole Z-pole run. Here, we quantify the triphoton branching fractions from meson decays by combining the exclusive $\rm Z\to\gamma+M$ results for spin-0,2 mesons of Tables 1113 with their corresponding two-photon branching fractions [ParticleDataGroup:2024cfk]. The results are listed in Table 15. The sum of all exclusive mesonic decay channels amounts to $\mathcal{B}=1.8\cdot 10^{-10}$, representing an increase of about 30% from the direct $3\gamma$ decay. The yields of SM triphoton decays will therefore amount to about 1 000 events for the $N(\rm Z)=6\cdot 10^{12}$ bosons expected to be collected during the full Z-pole operation at the FCC-ee. Searches for ALPs and new spin-2 particles will have to be carried out on top of such a relatively large number of background events that have been neglected so far in the derivation of ALP limits at future $\mathrm{e}^{+}\mathrm{e}^{-}$ facilities [Bauer:2018uxu].
@@
Table 15: Compilation of exclusive Z decays into triphoton final states. The predictions of relevant $\rm Z\to\gamma+M$ radiative decays listed in Tables 1113 are combined with the diphoton branching fractions of the produced mesons, $\rm M\to\gamma\gamma$ [ParticleDataGroup:2024cfk, Wang:2018rjg].
@@
$Z$ $\longrightarrow$ $\gamma\gamma\gamma$ $6.6$ $\times~10^{-10}$
@@ IV.2.4 Z boson decays into a W boson plus a meson
The exclusive Z boson decays into $\rm W^{\mp}$ boson plus a charged meson $\rm M^{\pm}$ correspond to the diagrams of Fig. 15 with W bosons in the final state. Theoretical estimations for such decays have been given in Ref. [Grossman:2015cak] for light-quark and $\rm D^{\pm}$ mesons. We estimate here the rest of missing exclusive decays: $\rm Z\to W^{\mp}+\{D^{*\pm},D^{*\pm}_{s},B^{\pm},B^{*\pm},B^{*\pm}_{c},B^{\pm}_{c}\}$, using the expression
@@
$\rm Z\,\to\,W$ ++ $\rm M$ Branching fraction Framework 2024 HL-LHC FCC-$e$
[…]
$\rm K^{\pm}$ $(1.16\pm 0.01)$ $\times~10^{-11}$ SCET+LCDA [Grossman:2015cak]
$\rm K^{*\pm}$ $(1.96\pm 0.12)$ $\times~10^{-11}$ SCET+LCDA [Grossman:2015cak]
$\rm D^{\pm}$ $(1.99\pm 0.17)$ $\times~10^{-11}$ SCET+LCDA [Grossman:2015cak]
$\rm D_{s}^{\pm}$ $(6.04\pm 0.30)$ $\times~10^{-10}$ SCET+LCDA [Grossman:2015cak]
@@
Figure 19: Branching fractions (in negative log scale) of exclusive $\rm Z\to W^{\pm}+meson$ decays. The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-$e$ reach based only on the total number of Z bosons to be produced.
@@ IV.3 Radiative Z boson leptonium decays
Figure 15 shows the diagrams for the decay of the Z boson into a photon plus a leptonium state. Depending on the relative helicities of the outgoing leptons, their bound state can be in the spin singlet (paraleptonium, $(\ell\ell)_{0}$) or triplet (ortholeptonium, $(\ell\ell)_{1}$) states. No calculation for such processes exists to our knowledge, but the key ingredients can be obtained from the determination of the $\rm\mathrm{e}^{+}\mathrm{e}^{-}\to\gamma+(\tau^{+}\tau^{-})_{0}$ cross section at the Z pole provided in the ditauonium studies of Ref. [dEnterria:2023yao]. The motivations for their study are the same as for the similar Higgs decays, namely, the search for LFV or LFUV phenomena.
@@
for the ortho-leptonium one, where $n$ is the principal quantum number of the leptonium resonance, and $s_{w}\equiv\sin\theta_{\mathrm{W}}$, and $c_{w}\equiv\cos\theta_{\mathrm{W}}$ are the sine and cosine of the Weinberg angle (Table 2). In both expressions above, $\alpha(0)$ and $\alpha(m_{\mathrm{Z}})$ are the electromagnetic coupling evaluated at zero [^12] The leptonium wavefunction is proportional to $\alpha(0)^{3}$, Eq. (24), which combined with the emission of a final-state onshell photon, gives the $\alpha(0)^{4}$ dependence shown in Eqs. (42) and (43). and at the Z pole mass, respectively, that, together with the rest of parameters, are listed in Table 2. The obtained branching fractions are tabulated in Table 17, and plotted in Fig. 21. The dependence of these Z-boson radiative rates on $\alpha^{6}\,m_{\ell\ell}^{2}$, leads to vanishingly small branching fractions, $\mathcal{O}(10^{-13}$$10^{-23})$, with ditauonium (positronium) featuring the largest (smallest) values. No experimental search has been conducted to date, although the relatively long lifetime of the leptonium objects (significantly boosted in the decay of the much heavier Z boson) would lead to a clean signature characterized by a displaced vertex from secondary decays of the $(\ell^{+}\ell^{-})$ into photons, $\rm e^{\pm}$, and/or $\mu^{\pm}$ [dEnterria:2022alo], akin to many BSM long-lived particles (but with an invariant mass at the corresponding known leptonium mass). Given the negligible rates, only an FCC-ee run at the Z pole would be able to provide limits approaching the SM values for the $\rm Z\to\gamma+(\tau^{+}\tau^{-})_{1}$ case.
@@
[…]
$\rm Z\,\to\,\gamma$ + $\rm(ee)_{1}$ $7.3$ $\times~10^{-21}$ Eq. (43) (this work)
$\rm(\mu\mu)_{1}$ $3.1$ $\times~10^{-16}$ Eq. (43) (this work)
$\rm(\tau\tau)_{1}$ $8.9$ $\times~10^{-14}$ Eq. (43) (this work)
$\rm(ee)_{0}$ $4.7$ $\times~10^{-23}$ Eq. (42) (this work)
$\rm(\mu\mu)_{0}$ $2.0$ $\times~10^{-18}$ Eq. (42) (this work)
$\rm(\tau\tau)_{0}$ $5.7$ $\times~10^{-16}$ Eq. (42) (this work)
Figure 21: Branching ratios (in negative log scale) of exclusive $\rm Z\to\gamma+leptonium$ decays. Theoretical predictions computed here are shown as blue bars. The red vertical line indicates the expected FCC-$e$ reach based on the total number of Z bosons to be produced.
@@ IV.4 Exclusive Z boson decays into two mesons
Figure 22 displays the diagrams of exclusive Z boson decays into two mesons, with contributions from direct quark decays (left and center), as well as from indirect $\rm V^{*}\to M$ transitions (right). Due to the coincidence of diagrams with those of the Higgs boson (Fig. 11), and since the Z boson yields at colliders are $10^{3}$$10^{6}$ times larger (Table 1), the study of such processes provides valuable information for the corresponding Higgs rates. The two-body decays into the same pair of pseudoscalar mesons, $\rm Z\to M+M$, are quantum-mechanically forbidden due to the presence of two identical final-state particles and conservation of angular momentum (thus violating the spin-statistics theorem), but the mesons M can be in a vector state. The decays of Z bosons into double quarkonia, first studied in 1990 [Bergstrom:1990bu], have been investigated as a means to provide clean information on the quarkonium bound-state dynamics at large momentum transfers. As explained below, there is still a relatively large uncertainty in the theoretical predictions, but this is still the most promising place to search for exclusive double-charmonia/bottomonia decays.
@@
To our knowledge, there are no calculations of exclusive decays of the Z boson involving light vector mesons, but only into charmonium and/or bottomonium states. We have evaluated here for the first time the double light vector meson decay $\rm Z\to VM+VM$ via the two direct (left and center) diagrams of Fig. 22, using the EFT + NRQM formalism of [Gao:2022mwa] in which the width can be written as
@@
for $N_{\mathrm{c}}=3$ colours, and without higher-order QCD corrections. We can only estimate the rate for the $\rm Z\to\phi+\phi$ channel because it is the only pure $|\mathrm{q}\overline{\mathrm{q}}\rangle$ state among the light mesons. Using Eqs. (44)–(47) and the parameters of Table 2, we obtain a $\mathcal{B}=2.1\cdot 10^{-12}$ rate (Table 18). There is no $\rm Z\to\phi+\phi$ experimental search performed to date, and only a machine like FCC-ee can provide enough Z bosons to start producing the decay. The table also quotes the forbidden $\rm Z\to\pi^{0}+\pi^{0}$ channel, for which the CDF upper bound could be improved by a factor of 100 at the HL-LHC. To provide a reasonable evaluation for exclusive double decays into other light mesons, rotation from flavour eigenstates $|\mathrm{u}\overline{\mathrm{u}}\rangle,|\mathrm{d}\overline{\mathrm{d}}\rangle$ to physical eigenstates such as $\rho,\omega,\eta$... would need to be performed. One can, however, anticipate branching fractions of the same order as the double-$\phi$ channel, $\mathcal{O}(10^{-12})$, since they only differ by a rotation and are enhanced by the second and third diagram of Fig. 22 involving extra $\rm\gamma^{*},W^{*}$ contributions. The scenario where Z decays into light pseudoscalar plus vector mesons suffers from the same complications of mixed states, but their rates are expected to be of the same order-of-magnitude as the double vector meson case because they have the same extra contribution from the photon propagator (Fig. 22, center) [Gao:2022mwa].
@@
$Z\,\to$ M ++ M Branching fraction Framework 2024 HL-LHC at FCC-ee
[…]
$\rm\pi^{+}$ + $\rm\pi^{-}$ $1.7$ $\times~10^{-12}$ SCET+LCDA [Cheng:2018khi]
$\rm K^{+}$ + $\rm K^{-}$ $8.3$ $\times~10^{-13}$ SCET+LCDA [Cheng:2018khi]
@@
There is a recent calculation of exclusive Z decays into a pair of charged pseudoscalar mesons, $\rm Z\to\pi^{+}\pi^{-},K^{+}K^{-}$ [Cheng:2018khi]. Also, the branching fractions $\rm Z\to B_{c}^{(*)+}+B_{c}^{(*)-}$ can be estimated to be in the $\mathcal{O}(10^{-8})-\mathcal{O}(10^{-11})$ range [Berezhnoy:2016etd, Liao:2024sgv, Liao:2022tqx], where the variation of results within a factor of 20 depends on the different potential models [Liao:2024sgv] or initial quark masses [Liao:2022tqx] used.
Tables 19 and 20 list the theoretical predictions and experimental limits for concrete two-meson decay modes with pairs of charmonium mesons and bottomonium plus other mesons, respectively. The same results are shown in graphical form in Figs. 23 and 24. The decay rates are very small, $\mathcal{B}(\rm Z\to M+M)\lesssim 10^{-10}$, and quite sensitive to the model details and ingredients. All heavy-quarkonium channels have rates obtained within the NRQCD + LDMEs formalism. Alternative LC calculations show a good agreement with the latter [Likhoded:2017jmx] but, unfortunately, they can only provide predictions for final states with different mesons or with mesons in different excited states. The NRQCD predictions from the work [Likhoded:2017jmx] show discrepancies with other later works, because they did not consider diagrams involving a photon propagator (Fig. 22, center) that can enhance the rates by more than two orders-of-magnitude. In contrast, this latter study provides results for some excited states that have not been calculated elsewhere (although their results for channels such as $\rm Z\to\chi_{c1}+\chi_{c1},\,\chi_{c2}+\chi_{c2},\,\chi_{c0}+\chi_{c2}$,…, which cannot be predicted using LC formalism, may need to be revisited for the same reasons stated above). In addition, recent calculations including higher-order QED [Luo:2022ugd] and QCD + QED [Li:2023tzx] corrections yield significant enhancements in the branching ratios for double-quarkonia production. All that said, and within the relatively large theoretical uncertainties for those processes, the largest branching fractions expected are for double-charmonium with $\mathcal{O}(10^{-10})$ rates. Despite a few searches carried out at LEP (and one at the LHC), experimental observation will only be possible at FCC-ee for about half of them.
@@
$Z\,\to$ M ++ M Branching fraction Framework 2024 HL-LHC at FCC-ee
$\rm Z\,\to$ $\rm\eta_{c}$ + $\rm J/\psi$ $(1.5\pm 0.4)$ $\times~10^{-11}$ NRQCD/NRCSM [Gao:2022mwa]
$(1.8\text{ -- }2.7)$ $\times~10^{-11}$ NRQCD+LDME (NLO) [Luo:2022ugd]
$2.7$ $\times~10^{-14}$ NRQCD+LDME [Likhoded:2017jmx]
$\rm\chi_{c0}$ $2.3$ $\times~10^{-12}$ NRQCD+LDME [Likhoded:2017jmx]
$(2.3\pm 1.0)$ $\times~10^{-12}$ LC+LCDA [Likhoded:2017jmx]
$\rm\chi_{c1}$ $5.4$ $\times~10^{-14}$ NRQCD+LDME [Likhoded:2017jmx]
$\rm h_{c}$ $2.1$ $\times~10^{-13}$ NRQCD+LDME [Likhoded:2017jmx]
$(1.0\pm 0.5)$ $\times~10^{-12}$ LC+LCDA [Likhoded:2017jmx]
$\rm\chi_{c2}$ $9.7$ $\times~10^{-13}$ NRQCD+LDME [Likhoded:2017jmx]
$(4.6\pm 2.0)$ $\times~10^{-12}$ LC+LCDA [Likhoded:2017jmx]
@@
$\left(1.11^{+0.34}_{-0.24}\right)$ $\times~10^{-10}$ NRQCD+LDME [Li:2023tzx]
(1.1 – 1.3)($1\text{ -- }$1.3) $\times~10^{-10}$ NRQCD+LDME (NLO) [Luo:2022ugd]
$2.3$ $\times~10^{-14}$ NRQCD+LDME [Likhoded:2017jmx]
$2.7$ $\times~10^{-11}$ NRQCD [Bergstrom:1990bu]
$\rm\chi_{c0}$ (1.1 – 4.1)($1\text{ -- }$4.1) $\times~10^{-12}$ NRQCD+LDME (NLO) [Luo:2022ugd]
$8.3$ $\times~10^{-14}$ NRQCD+LDME [Likhoded:2017jmx]
$(4.7\pm 2.0)$ $\times~10^{-13}$ LC+LCDA [Likhoded:2017jmx]
$\rm\chi_{c1}$ $(3.5\text{ -- }4.4)$ $\times~10^{-12}$ NRQCD+LDME (NLO) [Luo:2022ugd]
$3.5$ $\times~10^{-15}$ NRQCD+LDME [Likhoded:2017jmx]
$\rm h_{c}$ $1.5$ $\times~10^{-12}$ NRQCD+LDME [Likhoded:2017jmx]
$(9.5\pm 5.0)$ $\times~10^{-12}$ LC+LCDA [Likhoded:2017jmx]
$\rm\chi_{c2}$ $(9.6\text{ -- }24.8)$ $\times~10^{-13}$ NRQCD+LDME (NLO) [Luo:2022ugd]
$1.4$ $\times~10^{-13}$ NRQCD+LDME [Likhoded:2017jmx]
$(9.3\pm 4.0)$ $\times~10^{-13}$ LC+LCDA [Likhoded:2017jmx]
$\rm\chi_{c0}$ + $\rm\chi_{c1}$ $7.6$ $\times~10^{-14}$ NRQCD+LDME [Likhoded:2017jmx]
$(1.4\pm 1.0)$ $\times~10^{-12}$ LC+LCDA [Likhoded:2017jmx]
$\rm h_{c}$ $3.5$ $\times~10^{-16}$ NRQCD+LDME [Likhoded:2017jmx]
$\rm\chi_{c2}$ $6.4$ $\times~10^{-15}$ NRQCD+LDME [Likhoded:2017jmx]
$\rm\chi_{c1}$ + $\rm\chi_{c1}$ $3.9$ $\times~10^{-16}$ NRQCD+LDME [Likhoded:2017jmx]
$\rm h_{c}$ $2.9$ $\times~10^{-14}$ NRQCD+LDME [Likhoded:2017jmx]
$(6.1\pm 5.0)$ $\times~10^{-13}$ LC+LCDA [Likhoded:2017jmx]
$\rm\chi_{c2}$ $1.3$ $\times~10^{-13}$ NRQCD+LDME [Likhoded:2017jmx]
$(2.8\pm 2.0)$ $\times~10^{-12}$ LC+LCDA [Likhoded:2017jmx]
$\rm h_{c}$ + $\rm h_{c}$ $9.9$ $\times~10^{-17}$ NRQCD+LDME [Likhoded:2017jmx]
$\rm\chi_{c2}$ $2.3$ $\times~10^{-16}$ NRQCD+LDME [Likhoded:2017jmx]
$\rm\chi_{c2}$ + $\rm\chi_{c2}$ $1.3$ $\times~10^{-16}$ NRQCD+LDME [Likhoded:2017jmx]
Figure 23: Branching fractions (in negative log scale) of exclusive Z boson decays into two light mesons (upper) or two charmonium mesons (lower). The most recent theoretical predictions (blue bars) are compared to current experimental upper limits (violet) and expected conservative HL-LHC bounds (orange). The red vertical line indicates the expected FCC-$e$ reach based only on the total number of Z bosons to be produced.
Table 20: Compilation of exclusive Z decays to one bottomonium meson plus another meson. For each decay, we provide the predicted branching fraction(s) and the theoretical approach used to compute it, as well as the current experimental upper limit and that estimated for HL-LHC. The last column indicates whether the decay can be produced at FCC-ee.
Exp. limits Producible at
$Z\,\to$ M ++ M Branching fraction Framework 2024 HL-LHC at FCC-ee
$\rm Z\,\to$ $\rm\Upsilon(1S)$ + $\rm J/\psi$ $4.6$ $\times~10^{-11}$ NRQCD [Bergstrom:1990bu]
$\rm\eta_{b}(1S)$ $(1.9\pm 0.2)$ $\times~10^{-11}$ NRQCD/NRCSM [Gao:2022mwa]